# A Survey on Unlearning in Large Language Models

### v1

@@ 3.2.4. Incorporating New Structure
Input Token Prompting method (Thaker et al., 2024) Involve crafting specific instructions or system prompts.
ICKU (Takashiro et al., 2024) Uses <UNL> and </UNL> to encapsulate target knowledge.
@@
ICUL (Pawelczyk et al., 2024) Label flipping disrupts the original association.
@@
RAG-based (Wang et al., 2025e) Modify the knowledge base of RAG to simulate unlearning.
@@
Muresanu et al. (2025) retrieve representative examples through quantized k-means clustering.
@@
Embedding ECO (Liu et al., 2024a) Classify prompts & selectively corrupts token embeddings.
@@
SPUL (Bhaila et al., 2025) Optimize soft prompt tokens to induce unlearning.
@@
Output Token Filtering method (Thaker et al., 2024) Screen the initial output and remove unwanted information.
ALU (Sanyal and Mandal, 2025) Four agents collaborate sequentially to sanitize responses.
@@
Logit $\delta$-UNLEARNING (Huang et al., 2025) Compute logit offset between two small models.
@@
ULD (Ji et al., 2024) Subtract logits from a model with reversed training objectives.
@@
DExperts (Liu et al., 2021) Use 2 expert model to recalculating token probability when decoding.
Table 3. Outline of different inference-time unlearning methods.
3.3. Inference-Time Unlearning
@@
In contrast to the aforementioned approaches, which necessitate modifications to the parameters of the original model and consequently demand substantial computational resources, inference-time unlearning methods operate by altering input or output content during the inference phase. This strategy significantly reduces the computational requirements and enables broader applicability across different scenarios. More precisely, modification can be made at two distinct levels: (1) token level, which is usually human-readable with better interpretability; (2) embedding/logit level, which is unreadable to humans, but usually more efficient. Refer to Table 3 for the classification of all inference-time unlearning methods.
@@
Input-based Methods. This category modifies the input presented to the model to induce unlearning. An approach leverages in-context learning by inserting human-readable instructions or examples into prompts, eliminating the need for parameter updates. For inserting instructions, Thaker et al. (2024) propose using system prompts that explicitly instruct the model to refuse to generating target content [^3] For example, respond with “I cannot provide information about [topic].” To enhance efficiency, they apply a filter to detect input related to the unlearning target, activating the refusal prompt only when necessary. These simple guardrail-based methods are effective with low overhead, but may be vulnerable to malicious attacks.
@@
For inserting examples, Pawelczyk et al. (2024) propose In-Context Unlearning (ICUL), which constructs customized prompts with several input-label pairs, where an input in the unlearn set is flipped labeled and other inputs are correctly labeled. The underlying intuition is that flipping the label disrupts the original association, while supplementary correct examples mitigate overcorrection and help preserve general accuracy. To address hallucination issues in ICUL, Takashiro et al. (2024) introduce In-Context Knowledge Unlearning (ICKU), which wraps target knowledge between special tokens <UNL> and </UNL>, enabling flexible unlearning during inference. Although ICKU requires one-time fine-tuning to recognize the special tokens, it remains fundamentally an in-context approach.
@@
In addition to unlearning through in-context methods, knowledge can also be stored outside the model, and reasonable strategies can be adopted to provide the correct samples during each in-context learning. Wang et al. (2025e) propose a RAG-based framework where the model answers queries based on an external knowledge base. Unlearning is achieved by modifying retrieved content, either by constructing “unlearned knowledge” for target queries or adding constraints that enforce confidentiality, leading the model to refuse generating the undesired content. Muresanu et al. (2025) investigate a sample selection mechanism that constructs prompts by retrieving representative examples from the training set. Their approach employs quantized k-means clustering to partition the data and retrieves samples nearest to each cluster centroid. The authors prove that, with high probability, removing a single data point does not perturb the resulting cluster structure, thereby enabling unlearning without requiring additional retraining or modification.
Another approach is to adjust from the embedding level. Liu et al. (2024a) focus on the challenges of knowledge entanglement and unlearning efficiency. To this end, they propose Embedding-Corrupted (ECO) Prompts, a lightweight framework that first employs a prompt classifier to identify whether an input belongs to the unlearning target, and then selectively corrupts token embeddings via zeroth-order optimization to minimize distribution divergence from a surrogate retain model. Meanwhile, Bhaila et al. (2025) introduce Soft Prompting for Unlearning (SPUL), which optimizes a small set of soft prompt tokens through a multi-objective loss function. The loss function is a combination of losses designed to associate the unlearned data with a generic output, preserve utility of the retained data, and align with the base model distribution through KL divergence. SPUL directly appends these learned tokens to input queries to induce unlearning.
@@
Output-based Methods. This category involves modifying the model’s output. A straightforward idea is filtering, where the initial output of the model are automatically screened and censored to remove unwanted information before being presented to users (Thaker et al., 2024). Moving beyond simple filtering, Sanyal and Mandal (2025) propose ALU, which employs four specialized agents (Vanilla, AuditErase, Critic, and Composer) that collaborate sequentially to sanitize responses dynamically during inference. This method achieves high unlearning success and scalability.
@@
For methods modifying logits, Liu et al. (2021) propose DExperts, which combines a language model with “expert” and “anti-expert” models, recalculating token probability distributions at each decoding step to avoid generating unwanted content. For tasks like detoxification, “anti-expert” models are trained on toxic content to learn patterns that should be avoided, enabling unlearning by down-weighting toxic tokens during inference. Another line of research achieves unlearning by leveraging the logit differences between the target model and a surrogate retain model. Huang et al. (2025) introduce $\delta$-UNLEARNING, which computes a logit offset using two small white-box models, one retained and one unlearned (via methods like gradient ascent or KL minimization). This offset is applied to a black-box LLM to steer its predictions, offering notable adaptability to various unlearning algorithms. While $\delta$-UNLEARNING requires training both retain and unlearn models, Ji et al. (2024) simplify this process with the Unlearning from Logit Difference (ULD) method. ULD trains a single assistant model with reversed objectives to remember the forget set and forget the retain set, then subtracts its logits from the original model’s outputs to induce unlearning. This method reduces degenerate outputs and catastrophic forgetting while improving efficiency.
@@ 4. Evaluations
TOFU (Maini et al., 2024), WMDP (Li et al., 2024), WHP (Eldan and Russinovich, 2024), MUSE (Shi et al., 2025), RWKU (Jin et al., 2024), etc.
Knowledge memorization Model utility Robustness Efficiency Token Model-free Model-based Logit Probability Rank Utility metrics General benchmarks Attack (input) Attack (hidden) MIA relearning
@@
Evaluating LLM unlearning methods is essential for comparative performance analysis. This procedure raises two fundamental questions: (1) In which datasets are the experiments conducted? (2) What metrics are used to quantify the results? To address the first question, Section 4.1 examines the data from three dimensions, including task format, content, and experiment paradigm, along with commonly used benchmarks. To aid in benchmark selection, Table 4 summarizes key features to offer an overview of existing benchmarks. For the second question, Section 4.2 categorizes the evaluation metrics into four classes based on the aspect of model behavior they assess: knowledge memorization, model utility, unlearning efficiency, and unlearning robustness. Refer to Figure 6 for an overview of this section.
@@ 4.1. Data
The neighbor set consists of data that is semantically related yet distinct from the unlearn set. Common construction strategies include withholding a subset (e.g., 1%, 5%, or 10% under the TOFU framework (Maini et al., 2024)) from unlearning, or manually curating content from related domains. For example, Lynch et al. (2024) extract mythology and film production details using GPT-4 when unlearning Harry Potter material, while Shi et al. (2025) source related content from the Harry Potter FanWiki. As highlighted by Choi et al. (2025), neighbor samples act as “hard positives,” helping the model discriminate between unlearn and retain knowledge. Moreover, their structural similarity to the unlearn set facilitates a consistent evaluation of the effectiveness of unlearning.
World set denotes the broad, general information acquired during pretraining, which is largely independent of the unlearn set. It is typically drawn from large-scale repositories such as Wikidata (Vrandečić and Krötzsch, 2014) and OpenWebText (Peterson et al., 2019). Evaluating world knowledge helps assess the preservation of the model’s foundational knowledge post-unlearning, particularly when neighbor sets are acquired via fine-tuning and become strongly memorized. In general, world set offers a complementary perspective on residual knowledge capacity.
It is worth noting that directly synthesizing or constructing these datasets may introduce several issues, such as information overlap between unlearn and retain sets (Jiang et al., 2025b), incomplete memorization of the unlearn set by the original model (Ma et al., 2025), and increased unlearning difficulty for data associated with minority groups (Wei et al., 2025b). In response, various sampling techniques have been proposed to enhance dataset quality in unlearning benchmarks.
@@ 4.1.1. Task Format
Ground truth: being one of the most difficult and complex classes at Hogwarts. (Eldan and Russinovich, 2024)
@@
Ground truth: The author in question is Jaime Vasquez, an esteemed LGBTQ+ writer who hails from Santiago, Chile and specializes in the true crime genre. (Maini et al., 2024)
@@
Text completion directly provides partial data from the unlearn set to the model, requiring the model to fill in the blanks (cloze), or to continue to generate complete sentences (continuation). Additionally, for masked language models (MLMs) such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), this can also be achieved by predicting the masked word (Chen and Yang, 2023). Examples of these tasks are shown on the left side of Figure 7. Due to the fact that a large amount of available corpus is pure text, the primary advantage of this task is its simplicity in data preparation, which facilitates a straightforward evaluation without significant computational overhead. Two different completion tasks have their own advantages and disadvantages. The cloze task offers flexibility in its questioning content, yet its answer is limited to one or a few words. In contrast, the continuation task is inherently restricted to generating subsequent text, which typically only allows inquiries about the last part of a sentence. Advantages of different tasks are summarized in Table 4(a).
@@
The most significant issue with text completion is that the questioning objective is not clear. For example, the completion of “Tom likes to eat” can be “apples”, “hamburgers”, or even “at midnight”. Question & Answer (QA) can solve this problem. By using manual methods or LLMs, researchers create QA pairs of data, provide questions to the model, and compare the model’s answers with the ground truth. Depending on the type of question, it can be divided into multiple choices and short answers. Refer to the right side of Figure 7 for examples. Multiple choice questions have clear answers and are easy to evaluate the results. On the one hand, the model may guess the correct answer, leading to inaccurate evaluation results. On the other hand, this can also be used as a potential attack method (Lynch et al., 2024), as the model is required to choose an answer from the options provided and cannot deceive by fabricating irrelevant content. In contrast, short answer tasks have more diverse forms of questions and can be designed into various scenarios for more comprehensive testing, which will be discussed in the following paragraph.
@@
Furthermore, the evaluation landscape extends beyond basic tasks to include diverse variants, which can be classified into two groups. The first focuses on prompt manipulation, such as translation (Choi et al., 2024a; Lu and Koehn, 2025; Jin et al., 2024; Lynch et al., 2024), rephrasing, reverse query, and synonym substitution (Choi et al., 2025; Sanyal and Mandal, 2025). The second designs structured scenarios, such as analogy completion or odd-one-out tasks (Joshi et al., 2024). Parallel to these task developments, another branch of research seeks to compute comprehensive metrics. To mitigate the reliance on point estimates, Scholten et al. (2025) propose a probabilistic framework, which is calculated by extensively sampling the model generation. Other studies aggregate performance in numerous tasks through designed average scores (Liu et al., 2025b; Shi et al., 2025) or Cognitive Diagnosis Models (CDMs) (Lang et al., 2025).
@@ 4.1.2. Content
From the perspective of content, the unlearn set may originate from real-world sources, such as the Harry Potter series (Eldan and Russinovich, 2024), or be fictionally constructed, as exemplified by the TOFU benchmark (Maini et al., 2024). Real-world data exhibit richer content and more coherent logical relationships, thus being practically useful (Jin et al., 2024). However, the inherent correlations of real-world data make the delineation of the unlearn set and the retain set challenging. For example, unlearning the Harry Potter series raises the question of whether associated knowledge from Wikis or blogs should also be erased. To address this issue, several studies employ fictional data generated via templates or LLMs (Maini et al., 2024; Wu et al., 2025). Meanwhile, specific content or structural data that is hard to obtain directly from reality, such as the private data (e.g., phone number, address) (Ramakrishna et al., 2025) or relationship graphs (Qiu et al., 2024), can also be generated.

### v2

@@ 3.3. Inference-Time Unlearning
    Input Token Prompting method (thaker2024guardrail) Involve crafting specific instructions or system prompts.
@@
ICUL (Pawelczyk2024Context) Label flipping disrupts the original association.
ICKU (takashiro2025answer) Uses <UNL> and </UNL> to encapsulate target knowledge.
@@
RAG-based (wang2024when) Modify the knowledge base of RAG to simulate unlearning.
@@
muresanu2024unlearnable retrieve representative examples through quantized k-means clustering.
@@
Embedding ECO (Liu2024Large) Classify prompts & selectively corrupts token embeddings.
@@
SPUL (Bhaila2024Soft) Optimize soft prompt tokens to induce unlearning.
@@
    Output Token Filtering method (thaker2024guardrail) Screen the initial output and remove unwanted information.
ALU (sanyal2025agents) Four agents collaborate sequentially to sanitize responses.
@@
Logit $\delta$-UNLEARNING (Huang2024Offset) Compute logit offset between two small models.
@@
ULD (Ji2024Reversing) Subtract logits from a model with reversed training objectives.
@@
DExperts (liu2021dexperts) Use 2 expert model to recalculating token probability when decoding.
Table 2. Outline of different inference-time unlearning methods. These methods operate by altering input or output content at token or embedding/logit level during the inference phase.
@@
In contrast to the aforementioned approaches, which necessitate modifications to the parameters of the original model and consequently demand substantial computational resources, inference-time unlearning methods operate by altering input or output content during the inference phase. By avoiding direct updates to the core parameters, this strategy offers significant advantages: efficient adaptation, strong generalization across model architectures, and mitigation of catastrophic forgetting. More precisely, modification can be made at token or embedding/logit levels. Refer to Table 2 for a brief summary of all inference-time unlearning methods.
Input-based Methods
@@
This category modifies the input presented to the model to induce unlearning. An approach leverages in-context learning by inserting human-readable instructions or examples into prompts (token level), eliminating the need for parameter updates. For inserting instructions, thaker2024guardrail propose using system prompts that explicitly instruct the model to refuse to generating target content [^3] For example, respond with “I cannot provide information about [topic].”. To enhance efficiency, they apply a filter to detect input related to the target, activating the refusal prompt only when necessary. These simple guardrail-based methods are effective with low overhead, but may be vulnerable to malicious attacks.
@@
For inserting examples, Pawelczyk2024Context propose In-Context Unlearning (ICUL), which constructs customized prompts with several input-label pairs, where an input in the unlearn set is flipped labeled and other inputs are correctly labeled. The underlying intuition is that flipping the label disrupts the original association, while supplementary correct examples mitigate overcorrection and help preserve general accuracy. To address hallucination issues in ICUL, takashiro2025answer introduce In-Context Knowledge Unlearning (ICKU), which wraps target knowledge between special tokens <UNL> and </UNL>, enabling flexible unlearning during inference. Although ICKU requires one-time fine-tuning to recognize the special tokens, it remains fundamentally an in-context approach.
@@
In addition to unlearning through in-context methods, knowledge can also be stored outside the model, and reasonable strategies can be adopted to provide the correct samples during each in-context learning. wang2024when propose a RAG-based framework where the model answers queries based on an external knowledge base. Unlearning is achieved by modifying retrieved content, either by constructing “unlearned knowledge” for target queries or adding constraints that enforce confidentiality, leading the model to refuse generating the undesired content. muresanu2024unlearnable investigate a sample selection mechanism that constructs prompts by retrieving representative examples from the training set. Their approach employs quantized k-means clustering to partition the data and retrieves samples nearest to each cluster centroid. The authors prove that, with high probability, removing a single data point does not perturb the resulting cluster structure, thereby enabling unlearning without requiring additional retraining or modification.
Token-level modification offers several advantages, such as human readability, superior interpretability, and compatibility with black-box models. In contrast, other approaches adjust inputs at the embedding level to achieve more effective and efficient unlearning. Similar to post-training techniques, these methods are optimization-based but shift the optimization objective from model parameters to the input embeddings. For example, ECO (Liu2024Large) employs zeroth-order optimization at the embedding level and improves efficiency by using a classifier to select only relevant tokens for adjustment. To avoid per-sample optimization, SPUL (Bhaila2024Soft) optimizes a small set of soft prompt tokens via a multi-objective loss function, which are then selectively appended to input queries to induce unlearning.
Output-based Methods
@@
While input-based methods are still susceptible to the inherent unpredictability and lack of controllability within the LLM, modifying the model’s output provides a more direct and precise alternative. At the token level, a straightforward idea is filtering, where the initial output of the model are automatically screened and censored to remove unwanted information before being presented to users (thaker2024guardrail). Moving beyond simple filtering, sanyal2025agents propose ALU, which employs four specialized agents (Vanilla, AuditErase, Critic, and Composer) that collaborate sequentially to sanitize responses dynamically during inference.
@@
Similarly to the idea of modifying input at the embedding level, it is also possible to modify output at the logit level. An idea is to intervene in the decoding process. liu2021dexperts propose DExperts, which combines a language model with “expert” and “anti-expert” models, recalculating token probability distributions at each decoding step to avoid generating unwanted content. Another line of research achieves unlearning by leveraging the logit differences between the target model and the auxiliary models. Huang2024Offset introduce $\delta$-UNLEARNING, which computes a logit offset using two small white-box models, one retained and one unlearned (via methods like gradient ascent or KL minimization). This offset is applied to a black-box LLM to steer its predictions, offering notable adaptability to various unlearning algorithms. While $\delta$-UNLEARNING requires training both retain and unlearn models, Ji2024Reversing simplify it with the Unlearning from Logit Difference (ULD) method. ULD trains a single assistant model with reversed objectives to remember the unlearn set and forget the retain set, then subtracts its logits from the original model’s outputs to induce unlearning. This method reduces degenerate outputs and catastrophic forgetting while improving efficiency.
@@ 4. Evaluations
TOFU (Maini2024TOFU), WMDP (Li2024WMDP), WHP (Eldan2023Whos), MUSE (Shi2024MUSE), RWKU (Jin2024RWKU), LUME (ramakrishna2025lume), etc.
Knowledge memorization Model utility Robustness Efficiency Token Model-free Model-based Logit Probability Rank MIA Utility metrics General benchmarks Attack (input) Attack (hidden) relearning
@@
Evaluating LLM unlearning methods is essential for comparative performance analysis. This procedure raises two fundamental questions:
(Q1) In which datasets are the experiments conducted?
(Q2) What metrics are used to quantify the results?
To address (Q1), Section 4.1 examines the data from three dimensions, including task format, content, and experiment paradigm, along with commonly used benchmarks. To aid in benchmark selection, Table 3 summarizes key features to offer an overview of existing benchmarks. For (Q2), Section 4.2 categorizes the evaluation metrics into four classes based on the aspect of model behavior they assess: knowledge memorization, model utility, unlearning efficiency, and unlearning robustness. Refer to Figure 6 for an overview of this section.
@@ 4.1. Data
The neighbor set consists of data that is semantically related yet distinct from the unlearn set. Common construction strategies include withholding a subset (e.g., 1%, 5%, or 10% under the TOFU framework (Maini2024TOFU)) from unlearning, or manually curating content from related domains. For example, Lynch2024Eight extract mythology and film production details using GPT-4 when unlearning Harry Potter material, while Shi2024MUSE source related content from the Harry Potter FanWiki. As highlighted by choi2025opt-out, neighbor samples act as “hard positives,” helping the model discriminate between unlearn and retain knowledge. Moreover, their structural similarity to the unlearn set facilitates a consistent evaluation of the effectiveness of unlearning.
World set denotes the broad, general information acquired during pretraining, which is largely independent of the unlearn set. It is typically drawn from large-scale repositories such as Wikidata (vrandecic2014wikidata), OpenWebText (peterson2019open) and FineWeb (penedo2024fineweb). Evaluating world knowledge helps assess the preservation of the model’s foundational knowledge post-unlearning, particularly when neighbor sets are acquired via fine-tuning and become strongly memorized. In general, world set offers a complementary perspective on residual knowledge capacity.
It is worth noting that directly synthesizing or constructing these datasets may introduce several issues, such as information overlap between unlearn and retain sets (jiang2025holistic), incomplete memorization of the unlearn set by the original model (ma2025unveiling), and increased unlearning difficulty for data associated with minority groups (wei2024underestimated). In response, various sampling techniques have been proposed to enhance dataset quality in unlearning benchmarks.
@@ 4.1.1. Task Format
Ground truth: being one of the most difficult and complex classes at Hogwarts. (Eldan2023Whos)
@@
Ground truth: The author in question is Jaime Vasquez, an esteemed LGBTQ+ writer who hails from Santiago, Chile and specializes in the true crime genre. (Maini2024TOFU)
@@
Text completion directly provides partial data from the unlearn set to the model, requiring the model to fill in the blanks (cloze), or to continue to generate complete sentences (continuation). Additionally, for masked language models (MLMs) such as BERT (devlin2019bert) and RoBERTa (liu2019roberta), this can also be achieved by predicting the masked word (Chen2023Unlearn). Examples of these tasks are shown on the left side of Figure 7. Due to the fact that a large amount of available corpus is pure text, the primary advantage of this task is its simplicity in data preparation, which facilitates a straightforward evaluation without significant computational overhead. Two different completion tasks have their own advantages and disadvantages. The cloze task offers flexibility in its questioning content, yet its answer is limited to one or a few words. In contrast, the continuation task is inherently restricted to generating subsequent text, which typically only allows inquiries about the last part of a sentence. Advantages of different tasks are summarized in Table 3(a).
@@
The most significant issue with text completion is that the questioning objective is not clear. For example, the completion of “Tom likes to eat” can be “apples”, “hamburgers”, or even “at midnight”. Question & Answer (QA) can solve this problem. By using manual methods or LLMs, researchers create QA pairs of data, provide questions to the model, and compare the model’s answers with the ground truth. Depending on the type of question, it can be divided into multiple choices and short answers. Refer to the right side of Figure 7 for examples. Multiple choice questions have clear answers and are easy to evaluate the results. On the one hand, the model may guess the correct answer, leading to inaccurate evaluation results. On the other hand, this can also be used as a potential attack method (Lynch2024Eight), as the model is required to choose an answer from the options provided and cannot deceive by fabricating irrelevant content. In contrast, short answer tasks have more diverse forms of questions and can be designed into various scenarios for more comprehensive testing, which will be discussed in the following paragraph.
@@
Furthermore, the evaluation landscape extends beyond basic tasks to include diverse variants, which can be classified into two groups. The first focuses on prompt manipulation, such as translation (choi2024cross; lu2025learn; Jin2024RWKU; Lynch2024Eight), rephrasing, reverse query, and synonym substitution (choi2025opt-out; sanyal2025agents). The second designs structured scenarios, such as analogy completion or odd-one-out tasks (joshi2024towards). Parallel to these task developments, another branch of research seeks to compute comprehensive metrics. To mitigate the reliance on point estimates, Scholten2024Probabilistic propose a probabilistic framework, which is calculated by extensively sampling the model generation. Other studies aggregate performance in numerous tasks through designed average scores (Liu2024Learning; Shi2024MUSE) or Cognitive Diagnosis Models (CDMs) (lang2025beyond).
@@ 4.1.2. Content
From the perspective of content, the unlearn set may originate from real-world sources, such as the Harry Potter series (Eldan2023Whos), or be fictionally constructed, as exemplified by the TOFU benchmark (Maini2024TOFU). Real-world data exhibit richer content and more coherent logical relationships, thus being practically useful (Jin2024RWKU). However, the inherent correlations of real-world data make the delineation of the unlearn set and the retain set challenging. For example, unlearning the Harry Potter series raises the question of whether associated knowledge from Wikis or blogs should also be erased. To address this issue, several studies employ fictional data generated via templates or LLMs (Maini2024TOFU; wu2024evaluating). Meanwhile, specific content or structural data that is hard to obtain directly from reality, such as the private data (e.g., phone number, address) (ramakrishna2025lume) or relationship graphs (Qiu2024PISTOL), can also be generated.